# Large Continual Instruction Assistant

### v3

@@ Appendix I Instruction Grouping Results
Origin Answer with the option’s letter from the given choices directly Answer the question using a single word or phrase Answer the object in the image using a single word or phrase Please provide the bounding box coordinate of the region this sentence describes -
Diverse Answer with the option’s letter from the given choices directly Capture the essence of your response in a single word or a concise phrase Provide a succinct response with a single word or phrase Express the object in the image in a single word or a short, descriptive phrase Respond to the question briefly, using only one word or a phrase Answer the question using a single word or phrase Condense your answer for each question into a single word or concise phrase Please provide the bounding box coordinate of the region this sentence describes
10Type Answer with the option’s letter from the given choices directly Select the correct answer from the given choices and respond with the letter of the chosen option Determine the correct option from the provided choices and reply with its corresponding letter Pick the correct answer from the listed options and provide the letter of the selected option Identify the correct choice from the options below and respond with the letter of the correct option From the given choices, choose the correct answer and respond with the letter of that choice Choose the right answer from the options and respond with its letter Select the correct answer from the provided options and reply with the letter associated with it From the given choices, select the correct answer and reply with the letter of the chosen option Identify the correct option from the choices provided and respond with the letter of the correct option From the given choices, pick the correct answer and respond by indicating the letter of the correct option Use a single word or a short phrase to respond to the question Provide your answer in one word or a short, descriptive phrase Summarize your response in one word or a concise phrase Answer using only one word or a short, descriptive phrase Answer the question with a single word or a brief, descriptive phrase Use one word or a concise phrase to respond to the question Capture the essence of your response in a single word or a concise phrase Capture the essence of your response in one word or a short phrase Answer the question with just one word or a brief phrase Provide your answer in the form of a single word or a brief phrase Respond to the question using a single word or a brief phrase Provide your answer in the form of a single word or a concise phrase Respond to the question with a single word or a short phrase Respond to the question with just one word or a brief phrase Answer the question using a single word or a concise phrase Respond to the question with a single word or a brief phrase Provide your response using only one word or a short phrase Respond to the question using only one word or a concise phrase Answer the question with one word or a short phrase Respond with one word or a short phrase Answer the question with a single word or a brief phrase Respond to the question using just one word or a concise phrase Respond with one word or a concise phrase Provide your answer in the form of one word or a brief phrase Answer the question using only one word or a concise phrase Provide your answer using just one word or a short phrase Use one word or a brief phrase to answer the question Reply to the question using one word or a concise phrase Use a single word or a short phrase to answer the question Answer with a single word or a short phrase Answer the question using a single word or phrase Use one word or a short phrase to respond to the question Answer the question succinctly with one word or a brief phrase Use a single word or phrase to answer the question Provide an answer using only one word or a brief phrase Provide your answer to the question using only one word or a brief phrase Answer the question with just one word or a short phrase Answer the question using just one word or a concise phrase Respond to the question using a single word or a concise phrase Answer the question using a single word or phrase Summarize the object in the image in a single word or a brief phrase Provide the object in the image using a single word or a brief phrase Give the object in the image in the form of a single word or a concise phrase Express the object in the image with one word or a short, descriptive phrase Identify the type of content in the image using one word or a concise phrase Respond to the object in the image with a single word or a short, descriptive phrase Describe the content of the image using one word or a concise phrase Express the object in the image in a single word or a short, descriptive phrase Use a single word or a short phrase to categorize the image content Classify the image content using only one word or a brief phrase Use one word or a short phrase to classify the content of the image Extract and provide the coordinates for the bounding box described in the sentence Please provide the bounding box coordinate of the region this sentence describes Give the bounding box coordinates as described in the sentence Extract and provide the bounding box coordinates based on the region described in the sentence Find and provide the bounding box coordinates for the region mentioned in the sentence Identify and provide the bounding box coordinates that match the description given in this sentence Determine and provide the bounding box coordinates based on the description in the sentence Provide the coordinates for the bounding box based on the region described in the sentence Provide the coordinates of the bounding box that correspond to the region described in the sentence Identify and provide the coordinates of the bounding box described in the sentence Identify and give the coordinates of the bounding box as described by the sentence -
Appendix P Algorithm
@@
KwOut: training parameters pool ${\{f_{trn}^{i}\}}_{i=1}^{n}$, instruction codebook $\{I^{i}\}_{i=1}^{n}$. Initialize: ${\{f_{trn}^{i}\}}_{i=1}^{n}$, $\{I^{i}\}_{i=1}^{n}$. For $d$ = 1, …, $D$ 1. Collect the instructions of the current task $I_{c}$, match with $I^{i}$ in codebook, and obtain the maximum cosine similarity $s$. If $s\geq\epsilon$ 2. Initialize the current training parameters $f_{trn}$ and EMA parameters $f_{trn}^{*}$ from the matching parameters. Else 2. Random initialize the current training parameters $f_{trn}$ and EMA parameters $f_{trn}^{*}$. For $epoch$ = 1 For $t$ = 1, …, $T$ 3. Draw a mini-batch $B$ = ${\{({x}_{i}^{t},{I}_{i}^{t},{y}_{i}^{t})\}}_{i=1}^{n_{t}}$. For$({x_{t}},{I_{t}},{y_{t}})$ in $B$ 4. Prepend the $f_{trn}$ into the $f_{lfm}$ and obtain prediction $\hat{y}_{t}=f_{lfm}([x_{t};I_{t}])$. 5. Calculate per batch loss $\mathcal{L}_{B}$ by accumulating $\mathcal{L}_{x}(y,\hat{y})$ and update $f_{trn}$ with optimizer. 6. Record $f_{trn}$ and the corresponding gradients $\mathcal{L}^{\prime}$ at iteration $t$. 7. Calculate EMA weight $\beta_{t}$ according to Eq.(45). 8. Update $f_{trn}^{*}$ by Eq.(1). 9. Clear $f_{trn}$ and $\mathcal{L}^{\prime}$ at iteration $t-1$. If $s\geq\epsilon$ 10. Cover matching parameters in the pool with $f_{trn}^{*}$ and update the codebook with $I_{c}$. Else 10. Append $f_{trn}^{*}$ into the pool and update the codebook with $I_{c}$.

### v5

@@ Appendix I Instruction Grouping Results
Origin
Answer with the option’s
letter from the given
choices directly
Answer the question
using a single
word or phrase
Answer the object in the
image using a single
word or phrase
Please provide the bounding
box coordinate of the region
this sentence describes
Diverse
Answer with the option’s
letter from the given
choices directly
Capture the essence of your
response in a single word
or a concise phrase
Provide a succinct
response with a single
word or phrase
Express the object in
the image in
a single word or a
short, descriptive phrase
Respond to the question
briefly, using only one
word or a phrase
Answer the question
using a single
word or phrase
Condense your answer for
each question into a single
word or concise phrase
Please provide the bounding
box coordinate of the region
this sentence describes
10Type
Answer with the option’s letter from the given choices directly
Select the correct answer from the given choices and respond with the letter of the chosen option
Determine the correct option from the provided choices and reply with its corresponding letter
Pick the correct answer from the listed options and provide the letter of the selected option
Identify the correct choice from the options below and respond with the letter of the correct option
From the given choices, choose the correct answer and respond with the letter of that choice
Choose the right answer from the options and respond with its letter
Select the correct answer from the provided options and reply with the letter associated with it
From the given choices, select the correct answer and reply with the letter of the chosen option
Identify the correct option from the choices provided and respond with the letter of the correct option
From the given choices, pick the correct answer and respond by indicating the letter of the correct option
Use a single word or a short phrase to respond to the question
Provide your answer in one word or a short, descriptive phrase
Summarize your response in one word or a concise phrase
Answer using only one word or a short, descriptive phrase
Answer the question with a single word or a brief, descriptive phrase
Use one word or a concise phrase to respond to the question
Capture the essence of your response in a single word or a concise phrase
Capture the essence of your response in one word or a short phrase
Answer the question with just one word or a brief phrase
Provide your answer in the form of a single word or a brief phrase
Respond to the question using a single word or a brief phrase
Provide your answer in the form of a single word or a concise phrase
Respond to the question with a single word or a short phrase
Respond to the question with just one word or a brief phrase
Answer the question using a single word or a concise phrase
Respond to the question with a single word or a brief phrase
Provide your response using only one word or a short phrase
Respond to the question using only one word or a concise phrase
Answer the question with one word or a short phrase
Respond with one word or a short phrase
Answer the question with a single word or a brief phrase
Respond to the question using just one word or a concise phrase
Respond with one word or a concise phrase
Provide your answer in the form of one word or a brief phrase
Answer the question using only one word or a concise phrase
Provide your answer using just one word or a short phrase
Use one word or a brief phrase to answer the question
Reply to the question using one word or a concise phrase
Use a single word or a short phrase to answer the question
Answer with a single word or a short phrase
Answer the question using a single word or phrase
Use one word or a short phrase to respond to the question
Answer the question succinctly with one word or a brief phrase
Use a single word or phrase to answer the question
Provide an answer using only one word or a brief phrase
Provide your answer to the question using only one word or a brief phrase
Answer the question with just one word or a short phrase
Answer the question using just one word or a concise phrase
Respond to the question using a single word or a concise phrase
Answer the question using a single word or phrase
Summarize the object in the image in a single word or a brief phrase
Provide the object in the image using a single word or a brief phrase
Give the object in the image in the form of a single word or a concise phrase
Express the object in the image with one word or a short, descriptive phrase
Identify the type of content in the image using one word or a concise phrase
Respond to the object in the image with a single word or a short, descriptive phrase
Describe the content of the image using one word or a concise phrase
Express the object in the image in a single word or a short, descriptive phrase
Use a single word or a short phrase to categorize the image content
Classify the image content using only one word or a brief phrase
Use one word or a short phrase to classify the content of the image
Extract and provide the coordinates for the bounding box described in the sentence
Please provide the bounding box coordinate of the region this sentence describes
Give the bounding box coordinates as described in the sentence
Extract and provide the bounding box coordinates based on the region described in the sentence
Find and provide the bounding box coordinates for the region mentioned in the sentence
Identify and provide the bounding box coordinates that match the description given in this sentence
Determine and provide the bounding box coordinates based on the description in the sentence
Provide the coordinates for the bounding box based on the region described in the sentence
Provide the coordinates of the bounding box that correspond to the region described in the sentence
Identify and provide the coordinates of the bounding box described in the sentence
Identify and give the coordinates of the bounding box as described by the sentence
Appendix Q Algorithm
@@
KwOut: training parameters pool ${\{f_{trn}^{i}\}}_{i=1}^{n}$, instruction codebook $\{I^{i}\}_{i=1}^{n}$.
Initialize: ${\{f_{trn}^{i}\}}_{i=1}^{n}$, $\{I^{i}\}_{i=1}^{n}$.
For $d$ = 1, …, $D$
1. Collect the instructions of the current task $I_{c}$, match with $I^{i}$ in codebook, and obtain the maximum cosine similarity $s$.
If $s\geq\epsilon$
2. Initialize the current training parameters $f_{trn}$ and EMA parameters $f_{trn}^{*}$ from the matching parameters.
Else
2. Random initialize the current training parameters $f_{trn}$ and EMA parameters $f_{trn}^{*}$.
For epochepoch = 1
For $t$ = 1, …, $T$
3. Draw a mini-batch $B$ = ${\{({x}_{i}^{t},{I}_{i}^{t},{y}_{i}^{t})\}}_{i=1}^{n_{t}}$.
[…]
4. Prepend the $f_{trn}$ into the $f_{lfm}$ and obtain prediction $\hat{y}_{t}=f_{lfm}([x_{t};I_{t}])$.
[…]
6. Record $f_{trn}$ and the corresponding gradients $\mathcal{L}^{\prime}$ at iteration $t$.
7. Calculate EMA weight $\beta_{t}$ according to Eq.(45).
8. Update $f_{trn}^{*}$ by Eq.(1).
9. Clear $f_{trn}$ and $\mathcal{L}^{\prime}$ at iteration t1t-1.
If $s\geq\epsilon$
10. Cover matching parameters in the pool with $f_{trn}^{*}$ and update the codebook with $I_{c}$.
Else
10. Append $f_{trn}^{*}$ into the pool and update the codebook with $I_{c}$.